a trainer is teaching a dolphin to do tricks, the probability that the dolphin successfully performs tricks is 45%. out of 3 attempts, you want to find the probability that the dolphin fails exactly one time.
Answer #1

solution

Given that ,

p = 0.45

1 - p = 1 - 0.45 = 0.55

n =3

Using binomial probability formula ,

P(X = x) = (n C x) * px * (1 - p)n - x

P(X =1 ) = (3 C 1) * 0.451 * (0.55)2

probability =0.4084